how much of a 32% solution must be mixed with 48ml of a 40% solution to obtain a 35% solution? ______mL
Answer by checkley77(12844)

.32X+.40*48=.35(X+48)
.32X+19.2=.35X+16.8
.32X-.35X=16.8-19.2
-.03X-2.4
X=-2.4/.03
X=80 ML OF 32% SOLUTION IS NEEDED.
PROOF:
.32*80+.40*48=.35(80+48)
25.6+19.2=.35*128
44.8=44.8